OpenStudy (anonymous):

can someone show me the steps to solve this equation? Given the function f(x)=0.3(4)^x, what is the value of f^-1(6) @campbell_st

OpenStudy (campbell_st):

well you need to find the inverse of the function 1st so swap x and y and you have \[x = 0.3 \times 4^y\] now make y the subject of the equation...

OpenStudy (anonymous):

ok so would you square the equation?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

nope you wouldn't square the equation... the 1st step is to divide both sides of the equation by 0.3 \[\frac{x}{0.3} = 4^y\] whats the opposite of an exponent..?

OpenStudy (anonymous):

squaring the equation

OpenStudy (campbell_st):

nope.. sorry... that's not the opposite of an exponential

OpenStudy (anonymous):

hmmm would you multiply the 4 across the equation?

OpenStudy (campbell_st):

nope.... thats not it.... you need to find the opposite of an exponential

OpenStudy (anonymous):

oooh sorry, it would be something to do with log right?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

so it would be log4(x/3)=y?

OpenStudy (campbell_st):

thats right so \[\log(\frac{x}{0.3}) = y \times \log(4)\] so whats the next step

OpenStudy (anonymous):

ok hold on let me think about it

OpenStudy (campbell_st):

The only problem with taking a base 4 log is that you will need to use change of base at some point to evaluate the solution if your calculator only does base e or base 10 logs

OpenStudy (anonymous):

\[\log(x \div 3)/\log(4)=y\]

OpenStudy (anonymous):

is that looking correct @campbell_st

OpenStudy (campbell_st):

thats it... now substitute x = 6 and evaluate

OpenStudy (campbell_st):

except its 0.3.... not 3

OpenStudy (anonymous):

thank you :D @campbell_st

OpenStudy (anonymous):

ok the answer is 2.161 :DDDDD

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats what I got